# Generic Antiholomorphic Polymonial Vector Fields

### v1

@@ 5.2. Proof of the Realization Theorem
Let $h\colon\widetilde{M}\to\hat{\mathbb{C}}$ be a conformal mapping. Since we can post-compose by a Möbius transformation, we can suppose that $h(\infty)=\infty$. We can further normalize $h$ by requiring that $h$ in the chart $(\infty,\psi_{\infty})$ has the form $\zeta+o(\zeta)$. Note that $h$ is still not unique, as we can post-compost by an affine transformation $z\mapsto\lambda z+b$, where $\lambda$ is a $(k+2)$-root of unity. To fix the rotation, we require that the real interval $(R,\infty)$, for $R$ large enough, on $H_{1}$ be mapped by $h$ in the attracting sector at infinity that intersect $\mathbb{R}_{+}$.
@@ 6. Anti-polynomial vector fields with maximal number of heteroclinic connections
Given a fixed $(k+2)$-noncrossing tree, consider the family of structurally stable vector field $\overline{P_{\eta}}{\partial\partial z}$ in the generic stratum parametrized by the analytic invariant $\eta\in\mathbb{H}^{k}$. When $\Im\eta_{j}\to 0$, the corresponding sepal zone collapses to a curve. Indeed, in the rectifying coordinate, $\Im\eta_{j}$ is the width of the horizontal strip. If $\Re\eta_{j}>0$, then this curve contains a heteroclinic connection $\gamma$ and $\int_{\gamma}P=\Re\eta_{j}$. If $\Re\eta_{j}=0$, then the two simple singular points on the boundary of the sepal zone coalesce into a double saddle point. In both cases, the initial noncrossing tree is no longer adequate to describe the topology of the phase portrait.
@@
    Bifurcation of a multiple saddle point. Under perturbation in the right direction, a multiple points breaks in singular points of smaller codimension and gives rise to at least one sepal zone.
@@ 6.1. Case of $k$ heteroclinic connections
The topological invariant of this case has a beautiful simplicity: the separatrix graph, made of the union of heteroclinic connections, (outgoing and incoming) separatrices, and singular points, can be extended to a ternary tree by adding marked point at infinity as vertices. By ternary tree, we mean an ordered rooted tree for which each vertex is either a leaf or has three children. The root of the tree is the singular point attached to the outgoing separatrix tangent to $\mathbb{R}_{+}$ and identified with that marked point (so that the tree has $43k+$ 4 vertices and $33k+$ 3 vertices).
We say that a vertex is an internal vertex if it has children. We say that an edge is an internal edge if it is connected to two internal vertices. We define the following total order on the internal edges: we say that $e<f$ if either a path from the root to $f$ must first pass $e$ (see e.g. edges $1<2$ on Figure 10), or a direct path from the root to $f$ is further to right than a direct path from the root to $e$ (see e.g. edges $1<5$ or $4<6$ on  Figure 10).
@@
  • (Topological invariant) To every anti-polynomial vector field $\overline{P}{\partial\partial z}$ with $k$ heteroclinic connections, we can associate a unique ternary tree with $k+1$ internal vertices made with the singular points, the incoming and outgoing separatrices, the heteroclinic connections, and the marked points at infinity. The root of the tree is the singular point attached to outgoing separatrix tangent to $\mathbb{R}_{+}$.
@@
  • (Equivalence) Two such vector fields with the same topological invariant are quasi-conformally (and hence orbitally topologically) equivalent. Furthermore, if they also have the same analytic invariant, then they are equal.
  • (Realization) To every pair $(T_{k},\nu)$, where $T_{k}$ is a ternary tree with $k+1$ interior vertices and $\nu\in\mathbb{R}_{+}^{k}$, there exists a unique monic centred polynimial $P$ such that $\overline{P}{\partial\partial z}$ has $k$ heteroclinic connections and has $(T_{k},\nu)$ as invariants.
@@
Let $\mathcal{H}_{k}$ be the set of anti-polynomial vector fields of degree $k+1$ with $k$ heteroclinic connections. Then $\mathcal{H}_{k}/\sim_{\text{top}}$ has $A(k+2)={12(k+1)+1}\binom{3(k+1)}{k+1}$ equivalence classes, where $\sim_{\text{top}}$ is the topological orbital equivalence of vector fields.
@@
In Section 3.1, we noted that that those numbers are the same than the numbers we obtained for generic anti-holomorphic polynomial vector fields. Indeed, it is well known that the number of ternary tree with $k$ internal edges is $A(k+2)$. ∎
@@
The idea of the proof is the same as the Realization Theorem 5.1: we construct a Riemann surface isomorphic to $\hat{\mathbb{C}}$ with half-planes and push the vector field $1{\partial\partial Z}$ from the surface to $\hat{\mathbb{C}}$. This case is even simpler, since there are no horizontal strips, only half-planes. The heteroclinic connections will corresponds to intervals on the boundary of $2k$ of the half-planes.
First, we want to embed the ternary tree in $\overline{\mathbb{D}}$. We start by adding a vertex over the root and attaching it to the root with an edge; this corresponds to the marked point that was identified with the singular point. Now, we order the leafs from left to right, i.e. leaf 1 is the first leaf reached by a depth-first search in the tree by always going left, leaf 2 is the second leaf reached, and so on, and the added vertex over the root is leaf 0. Then we embed the tree in $\overline{\mathbb{D}}$ by mapping the $j$-th leaf on $e_{j}=\exp\!\left(2i\pi{j2k+2}\right)$ on $\partial\mathbb{D}$, and the internal vertices to points in $\mathbb{D}$ so that the edges can be mapped to simple curve that do not cross each other. See Figure 11.
@@
Figure 11. Example of a tree and its embeding.
@@
Each $H_{j}^{+}$ (resp. $H_{j}^{-}$) in $\mathbb{D}$ is adjacent to two edges and, say, $n_{j}$ internal edges. Suppose $\ell_{1}<\cdots<\ell_{n_{j}}$ are the internal edges. Then we mark $(-\infty,0)$ with $e_{2j-1}$, $(0,\nu_{\ell_{1}})$ by $\ell_{1}$, $\ldots$, $(\nu_{\ell_{1}}+\cdots+\nu_{\ell_{n_{j}-1}},\nu_{\ell_{1}}+\ldots+\nu_{\ell_{n_{j}}})$ by $\ell_{n_{j}}$, and lastly $(\nu_{\ell_{1}}+\cdots+\nu_{\ell_{n_{j}}},\infty)$ by $e_{2j-2}$ (resp. $e_{2j}$). We mark 00 and $\nu_{\ell_{1}},\ldots,\nu_{\ell_{n_{j}}}$ by the corresponding vertices of the tree.
@@ 6.2. Bifurcation diagram for quadratic anti-polynomial vector fields
Starting with $\arg\varepsilon=0$, the phase portrait has a heteroclinic connection oriented from $\sqrt{\varepsilon}$ to $-\sqrt{\varepsilon}$ (since $\int_{-\sqrt{\varepsilon}}^{\sqrt{\varepsilon}}P_{\varepsilon}<0$) on the real interval and the phase portrait is symmetric with respect to $\mathbb{R}$. When $\arg\varepsilon$ inscreases in the interval $(0,\pi)$, the singular point $\sqrt{\varepsilon}$ rotate counter-clocktwise around 0 until it reaches $i\mathbb{R}_{+}$. At $\arg\varepsilon={2\pi 3}$, there is again a heteroclinic connection oriented from $-\sqrt{\varepsilon}$ to $\sqrt{\varepsilon}$ (since $\int_{-\sqrt{\varepsilon}}^{\sqrt{\varepsilon}}P_{\varepsilon}>0$).
(a) Bifurcation diagram of $\dot{z}=\overline{z^{2}-\varepsilon}$.

### v2

@@ 5.2. Proof of the Realization Theorem
Let $h\colon\widetilde{M}\to\hat{\mathbb{C}}$ be a conformal mapping. Since we can post-compose by a Möbius transformation, we can suppose that $h(\infty)=\infty$. We can further normalize $h$ by requiring that $h$ in the chart $(\infty,\psi_{\infty})$ has the form $\zeta+o(\zeta)$. Note that $h$ is still not unique, as we can post-compose by an affine transformation $z\mapsto\lambda z+b$, where $\lambda$ is a $(k+2)$-root of unity. To fix the rotation, we require that the real interval $(R,\infty)$, for $R$ large enough, in $H_{1}$ be mapped by $h$ in the attracting sector at infinity that intersect $\mathbb{R}_{+}$.
@@ 6. Anti-polynomial vector fields with maximal number of heteroclinic connections
Given a fixed $(k+2)$-noncrossing tree, consider the family of structurally stable vector fields $\overline{P_{\eta}}{\partial\partial z}$ in the generic stratum parametrized by the analytic invariant $\eta\in\mathbb{H}^{k}$. When $\Im\eta_{j}\to 0$, the corresponding sepal zone collapses to a curve. Indeed, in the rectifying coordinate, $\Im\eta_{j}$ is the width of the horizontal strip. If $\Re\eta_{j}>0$, then this curve contains a heteroclinic connection $\gamma$ and $\int_{\gamma}P=\Re\eta_{j}$. If $\Re\eta_{j}=0$, then the two simple singular points on the boundary of the sepal zone coalesce into a double saddle point. In both cases, the initial noncrossing tree is no longer adequate to describe the topology of the phase portrait.
@@
    Bifurcation of a multiple saddle point. Under perturbation in the right direction, a multiple point breaks in singular points of smaller codimensions and gives rise to at least one sepal zone or heteroclinic connection.
@@ 6.1. Case of $k$ heteroclinic connections
The topological invariant of this case has a beautiful simplicity: the separatrix graph, made of the union of heteroclinic connections, (outgoing and incoming) separatrices, and singular points, can be extended to a ternary tree by adding marked points at infinity as vertices. By ternary tree, we mean an ordered rooted tree for which each vertex is either a leaf or has three children. In our case, we have one too many edge and vertex, so we collapse the edge corresponding to the separatrix tangent to $\mathbb{R}_{+}$ at infinity into a vertex, which we define to be the root. This way, the tree has $43k+$ 4 vertices and $33k+$ 3 edges.
We say that a vertex is an internal vertex if it has children. We say that an edge is an internal edge if it is connected to two internal vertices. We define the following total order on the internal edges: we say that $e<f$ if either a path from the root through $f$ must first pass $e$ (see e.g. edges $1<2$ on Figure 10), or a direct path from the root through $f$ is further to right than a direct path from the root through $e$ (see e.g. edges $1<5$ or $4<6$ on  Figure 10).
@@
  • (Topological invariant) To every anti-polynomial vector field $\overline{P}{\partial\partial z}$ with $k$ heteroclinic connections, we can associate a unique ternary tree with $k+1$ internal vertices made with the singular points, the incoming and outgoing separatrices, the heteroclinic connections, and the marked points at infinity. The root of the tree is the separatrix tangent to $\mathbb{R}_{+}$ collapsed into a vertex.
@@
  • (Equivalence) Two such vector fields with the same topological invariant are quasi-conformally (and hence orbitally topologically) equivalent. Furthermore, if they also have the same analytic invariant, then they are analytically conjugate.
  • (Realization) To every pair $(T_{k},\nu)$, where $T_{k}$ is a ternary tree with $k+1$ interior vertices and $\nu\in\mathbb{R}_{+}^{k}$, there exists a monic centred polynomial $P$ such that $\overline{P}{\partial\partial z}$ has $k$ heteroclinic connections and has $(T_{k},\nu)$ as invariants.
@@
Let $\mathcal{H}_{k}$ be the set of anti-polynomial vector fields of degree $k+1$ with $k$ heteroclinic connections. Then $\mathcal{H}_{k}/{\sim_{\text{top}}}$ has $A(k+2)={12(k+1)+1}\binom{3(k+1)}{k+1}$ equivalence classes, where $\sim_{\text{top}}$ is the topological orbital equivalence fixing marked points at infinity of vector fields .
@@
In Section 3.1, we noted that that those numbers are the same as the numbers we obtained for generic anti-holomorphic polynomial vector fields. Indeed, as discussed in that section, the number of ternary tree with $k$ internal edges is $A(k+2)$. ∎
@@
The idea of the proof is the same as the Realization Theorem 5.1: we construct a Riemann surface isomorphic to $\hat{\mathbb{C}}$ with half-planes and push the vector field $1{\partial\partial Z}$ from the surface to $\hat{\mathbb{C}}$. This case is even simpler, since there are no horizontal strips, only half-planes. The heteroclinic connections will correspond to intervals on the boundary of $2k$ of the half-planes.
First, we want to embed the ternary tree in $\overline{\mathbb{D}}$. We start by adding a vertex over the root and attaching it to the root with an edge; this corresponds to the outgoing separatrix tangent to $\mathbb{R}_{+}$ at infinity that was collapsed into a vertex. Now, we order the leafs from left to right, i.e. leaf 1 is the first leaf reached by a depth-first search in the tree by always going left, leaf 2 is the second leaf reached, and so on, and the added vertex over the root is leaf 0. Then we embed the tree in $\overline{\mathbb{D}}$ by mapping the $j$-th leaf on $e_{j}=\exp\!\left(2i\pi{j2k+2}\right)$ on $\partial\mathbb{D}$, and the internal vertices to points in $\mathbb{D}$ so that the edges can be mapped to simple curve that do not cross each other. See Figure 11.
@@
Figure 11. Example of a tree and its embedding.
@@
Each $H_{j}^{+}$ (resp. $H_{j}^{-}$) in $\mathbb{D}$ is adjacent to two edges and, say, $n_{j}^{\pm}$ internal edges. Suppose $\ell_{1}<\cdots<\ell_{n_{j}^{\pm}}$ are the internal edges. Then we mark $(-\infty,0)$ with $e_{2j-1}$, $(0,\nu_{\ell_{1}})$ by $\ell_{1}$, $\ldots$, $(\nu_{\ell_{1}}+\cdots+\nu_{\ell_{n_{j}^{\pm}-1}},\nu_{\ell_{1}}+\ldots+\nu_{\ell_{n_{j}^{\pm}}})$ by $\ell_{n_{j}^{\pm}}$, and lastly $(\nu_{\ell_{1}}+\cdots+\nu_{\ell_{n_{j}^{\pm}}},\infty)$ by $e_{2j-2}$ (resp. $e_{2j}$). We mark 00 and $\nu_{\ell_{1}},\ldots,\nu_{\ell_{n_{j}^{\pm}}}$ by the corresponding vertices of the tree.
@@ 6.2. Bifurcation diagram for quadratic anti-polynomial vector fields
Starting with $\arg\varepsilon=0$, the phase portrait has a heteroclinic connection oriented from $\sqrt{\varepsilon}$ to $-\sqrt{\varepsilon}$ (since $\int_{-\sqrt{\varepsilon}}^{\sqrt{\varepsilon}}P_{\varepsilon}<0$) on the real interval and the phase portrait is symmetric with respect to $\mathbb{R}$. When $\arg\varepsilon$ increases in the interval $(0,\pi)$, the singular point $\sqrt{\varepsilon}$ rotates counter-clockwise around 0 until it reaches $i\mathbb{R}_{+}$. At $\arg\varepsilon={2\pi 3}$, there is again a heteroclinic connection oriented from $-\sqrt{\varepsilon}$ to $\sqrt{\varepsilon}$ (since $\int_{-\sqrt{\varepsilon}}^{\sqrt{\varepsilon}}P_{\varepsilon}>0$).